### v2

@@ 3.1 A Tight Bound of $m/e$ on the Total Number of Copies
Fix $n$ MMS partitions $\{(P_{i}^{1},\dots,P_{i}^{n})\}_{i\in N}$. Let $A=(A_{1},\dots,A_{n})$ be a random allocation, where $A_{i}\sim U(\{P_{i}^{1},\dots,P_{i}^{n}\})$, independently for any agent $i$. That is, agent $i$ gets one of her MMS bundles chosen uniformly at random.
@@
where the third inequality follows from independence.
@@
3.2 Bound on the Maximum Number of Copies of Any Good
In this section we upper bound the number of copies of any individual good, and together with the bound established in Section 3.1, show to achieve both guarantees with high probability.
Theorem 2.
For any instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with monotone valuations, there exists a $(t,k)$-allocation that satisfies all agents and satisfies $t\leq m/e$ and $k\leq\min\left\{n-1,\frac{3\ln m}{\ln\ln m}\right\}$. Moreover, given an MMS partition for each agent, for any $\beta\in\mathbb{N}$, there exists an efficient algorithm that returns such an allocation with probability $1-e^{-\beta}$.
In the proof of Theorem 2, we use the following lemma. Its proof, which relies on a balls-and-bins argument, is deferred to Appendix B.
Lemma 2.
For any fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, with monotone valuations $\{v_{i}\}_{i\in N}$, there exists an MMS allocation where every good is copied at most $O\left(\frac{\ln m}{\ln\ln m}\right)$ times.
@@ 3.2 Bound on the Maximum Number of Copies of Any Good
Consider an allocation where agent $i$’s bundle, is randomly chosen as one of the sets in $i$’s MMS partition. We view this process as throwing $n$ balls (agents) into $n$ bins (the index in an MMS partition). Fix some good $g\in M$, it belongs to some MMS set for all the agents, possibly different indices, but we can rename the sets and assume without loss of generality that $g$ is always in the set of index 1 in the MMS partitions of all agents. The event in which $g$ is copied at least $k$ times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. For $k=\frac{3\ln m}{\ln\ln m}$, the probability of having a bin with more than $k$ balls is at most $\frac{1}{m}$. ∎
Proof of Theorem 2.
Fix an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$. For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be her MMS partition. Consider the algorithm $\mathcal{A}$ which does the following (at most) $3m\beta$ times:
@@
    Samples an allocation according to $A_{i}\sim U(P^{i})$.
@@
First, by the union bound, the probability that $\mathcal{A}$ fails in some iteration is
[…]
where the last inequality follows from the proof of Lemma 2.
Recall that in the proof of Proposition 1, we have established that when the allocation $A$ is picked uniformly at random, as in the above procedure, the expected number of total copies is $m/e$. This allows us to bound the second term using Markov’s inequality,
@@
[…]
4 MMS with Copies: Additive Valuations
In this section, we prove the following theorem.
Theorem 3.
For any instance with additive valuations, at most $\min\{n-2,\lfloor m/3\rfloor(1+o(1))\}$ distinct copies suffice to guarantee a full MMS allocation.
The proof of Theorem 3 is divided into two parts. In Section 4.1, we give an algorithm that obtains full MMS with $n-2$ distinct copies. In Section 4.2 we relate the $1$-out-of-$d$ bounds to our problem (a relation that holds for arbitrary monotone valuations), and use it to obtain full MMS with $\lfloor m/3\rfloor(1+o(1))$ distinct copies for additive valuations.
4.1 Full MMS with $n-2$ Copies
@@
       Let $P^{\prime}\subset P$ be the minimal set of bundles violating Hall’s theorem ($|P^{\prime}|>1$ since $1$ is matched to every node in $P$)
@@
       Find a matching between bundles in $\hat{P}$ and agents in $N$
@@
       Allocate agents in $\tilde{N}$ using procedure $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}((\tilde{N},\tilde{M},\{v_{i}\}_{i\in\tilde{N}}))$
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-2$ distinct copies.
Algorithm 1 $\mathsf{Match\mbox{-}n\mbox{-}Fill}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-2$ distinct copies.
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-1$ distinct copies.
Algorithm 2 $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-1$ distinct copies.
We first show that $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ satisfies all agents, assuming each agent values the remaining items high enough (proof in Appendix C). An example of a run of $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ appears in Appendix E.
Lemma 3.
Assuming that for every $i\in N$, $v_{i}(M)\geq|N|$ and that for every item $j$, $v_{ij}<1$, $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ outputs an allocation such that $v_{i}(A_{i})\geq 1$ for every $i\in N$, and creates at most $|N|-1$ distinct copies.
@@
A simple induction shows that every time we allocate a bundle, the value of every remaining agent $i$ for all remaining items is at least the number of remaining agents times $i$’s MMS value. Indeed, $i$’s value for all items except the last item inserted to the bag is at most $1$, otherwise, $i$ could have been allocated in a previous iteration. Since we duplicate the last inserted item, we get that the number of agents decreased by 1, while the total value of $i$ for all remaining items decreased by at most 1. Thus, we make sure that at any point, we have enough value to satisfy all remaining agents. ∎
We now show that $n-2$ distinct copies suffice.
Lemma 4.
For any fair division instance with additive valuations, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ guarantees a full MMS allocation for all the agents with at most $n-2$ distinct copies.
@@
First, by definition, every agent that is assigned some bundle for which they have an edge in $G$ gets their MMS value. Moreover, since agent $1$ has an edge to all bundles in $P$, then at least one agent gets satisfied, and $|\tilde{N}|\leq|N|-1$. It remains to show that sets $\tilde{N}$ and $\tilde{M}$ satisfy the conditions of Lemma 3.
$\mathsf{Match\mbox{-}n\mbox{-}Fill}$ first allocates items with value more than the MMS of an agent, so for every $i\in\tilde{N},j\in\tilde{M}$, $v_{ij}<1$. Moreover, by Lemma 1, allocating a single item to a single agent is a valid reduction. Thus, after this step, we have that for every $i\in N$, $v_{i}(M)\geq|N|$. We show that after allocating bundles $\hat{P}$ to agents in $N$, for every unallocated agent $i\in\tilde{N}$, $v_{i}(\tilde{M})\geq|\tilde{N}|$.
Recall that $P^{\prime}$ is a minimal set of bundles violating Hall’s Theorem, and let $|P^{\prime}|=k$. Let $N^{\prime}=\Gamma(P^{\prime})$ be their neighbors in $G$. Note that $|N^{\prime}|=k-1$, otherwise, one can remove an arbitrary element of $P^{\prime}$ and get a smaller set violating Hall’s Theorem. Moreover, since $\hat{P}$ does not violate Hall’s theorem, we know that $\Gamma(\hat{P})=N^{\prime}$ and that the agents in $N^{\prime}$ are exactly the agents matched by bundles from $\hat{P}$. Thus, it must be that there are no edges between agents in $\tilde{N}=N\setminus N^{\prime}$ and bundles in $\hat{P}$. Therefore, for every $i\in\tilde{N}$, $v_{i}(\bigcup_{S\in\hat{P}}S)\leq k-1$, which implies
$v_{i}(\tilde{M})=v_{i}(M)-v_{i}\left(\bigcup_{S\in\hat{P}}S\right)\geq|N|-(k-1)=|N|-|N^{\prime}|=|\tilde{N}|,$
as desired. It follows that all the conditions required in order to apply Lemma 3 are fulfilled, and we can satisfy agents in $\tilde{N}$ using $|\tilde{N}|-1\leq|N|-2\leq n-2$ distinct copies (recall that at least one agent was already allocated a bundle before). Since $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is the only place where we produce copies, this concludes the proof. ∎
Together with the example in [27] of an instance with $n=3$ additive agents that does not admit an MMS allocation, Lemma 4 implies a tight bound on the number of copies for $3$ additive agents.
Corollary 2.
For every instance with 3 additive valuations, a single copy of a single item suffices to guarantee a full MMS allocation. This is tight, i.e., there are instances where one copy is required.
Remark.
One can consider a similar question in the context of chore allocation, and ask “what is the maximum number of chores that can be assigned to agents such that no agent’s cost is larger than their maximin share”; that is, how many chores are needed to be set aside in order to produce a full MMS allocation. It turns out that an analogous process to $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ in the context of chores outputs an allocation where at most $n-2$ items are disposed of, and each agent is allocated a bundle of cost no larger than their MMS value. More on this is in Appendix A.
4.2 Full MMS with $\lfloor m/3\rfloor(1+o(1))$ Copies
@@
Lemma 5.
If there exists a 1-out-of-$(1+\alpha)n$ MMS allocation for every fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ for some constant $\alpha\in(0,1)$ then
@@
distinct copies suffice to guarantee a full MMS allocation. This holds for arbitrary monotone valuations.
@@
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance for which there exists a 1-out-of-$(1+\alpha)n$ MMS. Notice that if we take a set of agents $N^{\prime}$ of size $n^{\prime}$ for which $(1+\alpha)n^{\prime}\leq n$ then we can use the guarantee stated in the lemma to produce an allocation $\{A_{i}\}_{i\in N^{\prime}}$ without copies such that for each agent $i\in N^{\prime}$,
@@
Since we chose the set of $|\tilde{N}|$ agents with the least number of items in $N^{\prime}$, it holds that
@@
We apply the above lemma to the following state-of-the-art bound due to [4].
Lemma 6 (Akrami et al. [4]).
@@
Corollary 3.
For any fair division instance with additive valuations, at most $\lfloor\frac{m}{3}\rfloor(1+o(1))$ distinct copies suffice to guarantee a full MMS allocation.
@@
Since $4\lceil n/3\rceil\leq 4n/3+4=n(1+\frac{1}{3}+\frac{4}{n})$, applying Lemma 5 with $\alpha=\frac{1}{3}+\frac{4}{n}$ implies that the number of copies sufficient in order to produce a full MMS allocation is at most
@@
$=$ $\left\lfloor\frac{m}{3}\right\rfloor(1+o(1)).$
5 Approximate-MMS with Copies: Additive Valuations
In this section we study approximate MMS allocations while allowing duplication of goods.

### v3

@@ 3.1.2 Upper Bound on the Total Number of Copies for General Monotone Valuations
Fix $n$ MMS partitions $\{P^{i}=(P^{i}_{1},\dots,P^{i}_{n})\}_{i\in N}$. Let $A=(A_{1},\dots A_{n})$ be a random allocation, where $A_{i}\sim\mathrm{Unif}(\{P^{i}_{1},\dots,P^{i}_{n}\})$, independently for any agent $i$. That is, agent $i$ gets one of her MMS bundles chosen uniformly at random.
@@
where the third inequality follows from independence across the agents.
@@
3.1.3 Upper Bound of $\ln m/\ln\ln m$ on the Maximum Number of Individual Copies
In this section, we upper bound the number of copies of any individual good, and together with the bound established in Section 3.1.2, show how to achieve both guarantees with high probability.
Theorem 3.4.
For any instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ with monotone valuations, there exists a $(t,k)$-allocation that ensures MMS for all the agents and satisfies $t\leq m/e$ along with $k\leq\min\left\{n-1,\frac{3\ln m}{\ln\ln m}\right\}$. Moreover, given an MMS partition for each agent, for any $\beta\in\mathbb{N}$, there exists an efficient algorithm that returns such an allocation with probability at least $1-e^{-\beta}$.
Notably, the upper bounds on the total and individual number of copies established in Theorem 3.4 are tight by Theorem 3.1, where one good is copied $n-1\geq\frac{\ln m}{\ln\ln m}-1$ times. In the proof of Theorem 3.4, we use the following lemma. Its proof, which relies on a balls-and-bins argument, is deferred to Appendix B.
Lemma 3.5.
For any fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, with monotone valuations $\{v_{i}\}_{i\in N}$, there exists an MMS allocation where every good is copied at most $O\left(\frac{\ln m}{\ln\ln m}\right)$ times.
@@ 3.1.3 Upper Bound of $\ln m/\ln\ln m$ on the Maximum Number of Individual Copies
Consider an allocation where agent $i$’s bundle is a set from her MMS partition, picked uniformly at random. We view this process as throwing $n$ balls (agents) into $n$ bins (the index in an MMS partition). Fix some good $g\in M$, it belongs to some MMS bundle for all the agents, possibly different indices, but we can re-index the sets and assume without loss of generality that $g$ is always in the set of index 1 in the MMS partitions of all agents. The event in which $g$ is copied at least k1k-1 times for the allocation $A=(A_{1},\dots A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. For $k=\frac{3\ln m}{\ln\ln m}$, the probability of having a bin with more than $k$ balls is at most $\frac{1}{m}$. ∎
Proof of Theorem 3.4.
Fix an instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$. For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be her MMS partition. Consider the algorithm $\mathcal{A}$ which does the following (at most) $3m\beta$ times:
@@
    Samples an allocation according to $A_{i}\sim\mathrm{Unif}(P^{i})$.
@@
First, by the union bound, the probability that algorithm $\mathcal{A}$ fails in some iteration is
$\Pr[\mathcal{A}\text{ fails in a single iteration}]\leq\Pr[t>m/e]+\Pr\left[k>\frac{3\ln m}{\ln\ln m}\right]\leq\Pr[t>m/e]+\frac{1}{m},$
where the last inequality follows from the proof of Lemma 3.5.
Recall that in the proof of Proposition 3.3, we have established that when the allocation $A$ is picked uniformly at random, as in the above procedure, the expected number of total copies is $m/e$. This allows us to bound the second term using Markov’s inequality,
@@
$\Pr[\mathcal{A}\text{ fails in a single iteration}]$ $\leq\frac{m}{m+2}+\frac{1}{m}=\frac{m^{2}+m+2}{m^{2}+2m}\leq\frac{m+3/2}{m+2}\leq 1-\frac{1}{3m}.$
$\Pr[\mathcal{A}\text{ fails in all iterations}]$ $=\left(\Pr[\mathcal{A}\text{ fails}]\right)^{3m\beta}\leq\left(1-\frac{1}{3m}\right)^{3m\beta}\leq\frac{1}{e^{\beta}}.\qed$
3.2 Additive Valuations
In this section, we prove the following theorem.
Theorem 3.6.
For any instance with additive valuations, at most $\min\{n-2,\lfloor m/3\rfloor(1+o(1))\}$ distinct copies suffice to guarantee an exact MMS allocation.
The proof of Theorem 3.6 is divided into two parts. In Section 3.2.1, we give an algorithm that obtains exact MMS with n2n-2 distinct copies. In Section 3.2.2 we relate the 11-out-of-$d$ bounds to our problem (a relation that holds for arbitrary monotone valuations, and use it to obtain exact MMS with $\lfloor m/3\rfloor(1+o(1))$ distinct copies for additive valuations.
@@ 3.2 Additive Valuations
   Let $P^{\prime}\subset P$ be the minimal set of bundles violating Hall’s condition ($|P^{\prime}|>1$ since $1$ is matched to every node in $P$)
@@
   Find a matching between bundles in $\hat{P}$ and agents in NN
@@
   Allocate agents in $\tilde{N}$ using procedure $)\mathsf{BagFill\mbox{-}with\mbox{-}Copies}((\tilde{N},\tilde{M},\{v_{i}\}_{i\in\tilde{N}}$))
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most n2n-2 distinct copies.
Algorithm 1 $\mathsf{Match\mbox{-}n\mbox{-}Fill}(\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle)$:
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most n2n-2 distinct copies.
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most n1n-1 distinct copies.
Algorithm 2 $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}(\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle)$:
@@
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most n1n-1 distinct copies.
We first show that $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ satisfies all agents, assuming each agent values the remaining items high enough (proof in Appendix B). An example of a run of $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ appears in Section B.1.
Lemma 3.7.
Assuming that for every $i\in N$, $v_{i}(M)\geq|N|$ and that for every item $j$, $v_{ij}<1$, $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ outputs an allocation such that $v_{i}(A_{i})\geq 1$ for every $i\in N$, and creates at most |N|1|N|-1 distinct copies.
@@
A simple induction shows that every time we allocate a bundle, the value of every remaining agent $i$ for all remaining items is at least the number of remaining agents. Indeed, $i$’s value for all items except the last item inserted to the bag is at most $1$, otherwise, $i$ could have been allocated in a previous iteration. Since we duplicate the last inserted item, we get that the number of agents decreased by 1, while the total value of $i$ for all remaining items decreased by at most 1. Thus, we make sure that at any point, we have enough value to satisfy all remaining agents. ∎
We now show that n2n-2 distinct copies suffice.
Lemma 3.8.
For any fair division instance with additive valuations, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ guarantees an exact MMS allocation for all the agents with at most n2n-2 distinct copies.
@@
First, by definition, every agent that is assigned some bundle for which they have an edge in $G$ gets their MMS value. Moreover, since agent $1$ has an edge to all bundles in $P$, then at least one agent gets satisfied, and $|\tilde{N}|\leq|N|-1$. It remains to show that sets $\tilde{N}$ and $\tilde{M}$ satisfy the conditions of Lemma 3.7.
[…]
Recall that $P^{\prime}$ is a minimal set of bundles violating Hall’s condition, and let $|P^{\prime}|=k$. Let $N^{\prime}=\Gamma(P^{\prime})$ be their neighbors in $G$. Note that $|N^{\prime}|=k-1$, otherwise, one can remove an arbitrary element of $P^{\prime}$ and get a smaller set violating Hall’s condition. Moreover, since $\hat{P}$ does not violate Hall’s condition, we know that $\Gamma(\hat{P})=N^{\prime}$ and that the agents in $N^{\prime}$ are exactly the agents matched by bundles from $\hat{P}$. Thus, it must be that there are no edges between agents in $\tilde{N}=N\setminus N^{\prime}$ and bundles in $\hat{P}$. Therefore, for every $i\in\tilde{N}$, $v_{i}(\bigcup_{S\in\hat{P}}S)\leq k-1$, which implies
[…]
as desired. It follows that all the conditions required in order to apply Lemma 3.7 are fulfilled, and we can satisfy agents in $\tilde{N}$ using $|\tilde{N}|-1\leq|N|-2\leq n-2$ distinct copies (recall that at least one agent was already allocated a bundle before). Since $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is the only place where we produce copies, this concludes the proof. ∎
Together with the example in [feige2021tight] of an instance with $n=3$ additive agents that does not admit an MMS allocation, Lemma 3.8 implies a tight bound on the number of copies for $3$ additive agents.
Corollary 3.9.
For every instance with 3 additive agents, a single copy of a single item suffices to guarantee an exact MMS allocation. This is tight, i.e., there are instances where one copy is required.
3.2.2 Exact MMS with $\lfloor m/3\rfloor(1+o(1))$ Copies
@@
Lemma 3.10.
If there exists a 1-out-of-$(1+\alpha)n$ MMS allocation for every fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ for some constant $\alpha\in(0,1)$ then
@@
distinct copies suffice to guarantee an exact MMS allocation. This holds for arbitrary monotone valuations.
@@
Let $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ be a fair division instance for which there exists a 1-out-of-$(1+\alpha)n$ MMS. Notice that if we take a set of agents $N^{\prime}$ of size $n^{\prime}$ for which $(1+\alpha)n^{\prime}\leq n$ then we can use the guarantee stated in the lemma to produce an allocation $\{A_{i}\}_{i\in N^{\prime}}$ without copies such that for each agent $i\in N^{\prime}$,
@@
[…]
@@
We apply the above lemma to the following state-of-the-art bound due to [akrami2023improving].
Lemma 3.11 ([akrami2023improving]).
@@
Corollary 3.12.
For any fair division instance with additive valuations, at most $\lfloor\frac{m}{3}\rfloor(1+o(1))$ distinct copies suffice to guarantee an exact MMS allocation.
@@
Since $4\lceil n/3\rceil\leq 4n/3+4=n(1+\frac{1}{3}+\frac{4}{n})$, applying Lemma 3.10 with $\alpha=\frac{1}{3}+\frac{4}{n}$ implies that the number of copies sufficient in order to produce an exact MMS allocation is at most
@@
[…]
In this section we prove that there exists an instance where agents have submodular valuations and any MMS allocation requires at least n1n-1 total copies. Our construction uses ideas from the submodular instance of [ghodsi2022fair].
Consider $n$ agents and $n^{n}$ goods positioned on the $n$-dimensional cube, $M=\{g_{d_{1},\dots,d_{n}}\mid d_{i}\in[n]\;\;\forall i\in[n]\}$. For any agent $i$ let $P_{i}^{j}=\{g_{d_{1},\dots d_{n}}\mid d_{i}=j\}$, similar to the definition in Section 3.1.1. Agent $i$’s valuation, $v_{i}$, is such that the MMS partition of agent $i$ is $P_{i}=\{P_{i}^{j}\}_{j\in[n]}$.
[…]
Observation 3.13.
$v_{i}$ is submodular.
Proof.
The marginal value of all goods is 1 for $|S|<n^{n-1}-1$. Take $|S|=n^{n-1}-1$. If $S\cup\{g\}=P_{i}^{j}$, the marginal value of $g$ is 1 and no more than $1/2$ for any super set of $S$. If $S\cup\{g\}\neq P_{i}^{j}$, then the marginal value of $g$ is $1/2$, and no more than $1/2$ for any super set of $S$. ∎
Proposition 3.14.
There exists an instance with submodular valuations such that any allocation with copies that achieve MMS, must have at least n1n-1 total copies.
Proof.
Let $A$ be an allocation (with copies), which achieves MMS. Observe that $A$ must satisfy $v_{i}(A_{i})\geq n^{n-1}$ for any $i\in[n]$. If agent $i$ does not get one of her MMS bundles, then to satisfy the above, it must be that $|A_{i}|\geq n^{n-1}+1$. Additionally, by using less than $n^{n-2}$ copies, only one agent $i$ can get a bundle $P^{j}_{i}$ for some $j\in[n]$, as any two bundle $P_{i}^{j}$ and $P_{i^{\prime}}^{j^{\prime}}$ for $i\neq i^{\prime}$ share $n^{n-2}$ goods: those with $d_{i}=j$ and $d_{i^{\prime}}=j^{\prime}$. Thus, the total number of items in $A$, including copies is at least $\sum_{i=1}^{n}|A_{i}|\geq n^{n-1}+(n-1)\cdot(n^{n-1}+1)=n^{n}+n-1$. This concludes the proof. ∎
4 Approximate Maximin Shares
In this section we study approximate MMS allocations while allowing duplication of goods, assuming all agents have additive valuations.